# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v4

@@ 6.5. Latency Estimation and Scalability
Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024a, b), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time. We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops. Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented. The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds. Only the kernel 3mm with tree reduction timeout at 4h. 32/36 NLP’s problem are solve in less than 1 minutes. It is important to note that the solver provides a feasible solution early in the process, though it may not be theoretically optimal. As the search continues, the solution improves, and once the solver completes, the solution is theoretically optimal. Therefore, it is possible to set a timeout to obtain a feasible, but not optimal, solution more quickly.
@@ 8. Limitations
Sisyphus is designed to optimize affine loop nests. For applications with non-affine code, we can partition the code into affine and non-affine sections. Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts. The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming. This combined approach ensures that both affine and non-affine portions of the application are handled in an optimal manner, achieving a balance between different types of optimizations and seamlessly integrating them for improved overall performance.
@@ 9. Conclusion
In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through high-level synthesis. By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods. Sisyphus defines a design space in the form of a template. Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR. Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of QoR.

### v5

@@ 6.5. Latency Estimation and Scalability
Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024, 2025), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time. We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops. Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented. As noted in Section 5, our post-optimization focuses primarily on memory transfer overlap, so our final latency values do not directly reflect the model’s initial predictions. However, our model achieves (when predicting the latency of the final design generated) an average prediction error of 8% across evaluated designs, with a 3% error in geometric mean. For certain designs, like syrk and syr2k, where the compiler achieves an II=2 rather than the expected II=1, prediction errors are higher (49% and 32%, respectively) due to an optimistic estimation of II. The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds. Only the kernel 3mm with tree reduction timeout at 4h. 32/36 NLP’s problem are solve in less than 1 minutes. It is important to note that the solver provides a feasible solution early in the process, though it may not be theoretically optimal. As the search continues, the solution improves, and once the solver completes, the solution is theoretically optimal. Therefore, it is possible to set a timeout to obtain a feasible, but not optimal, solution more quickly.
@@ 8. Limitations
Sisyphus is designed to optimize affine loop nests. For applications with non-affine code, we can partition the code into affine and non-affine sections. Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2022b) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts. The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming. This combined approach ensures that both affine and non-affine portions of the application are handled in an optimal manner, achieving a balance between different types of optimizations and seamlessly integrating them for improved overall performance.
@@ 9. Conclusion
In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through HLS. By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods. Sisyphus defines a design space in the form of a template. Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR. Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of quality of result. Acknowledgments -   This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD [^1] J. Cong has a financial interest in AMD.
/HACC Program.